OpenStudy (anonymous):

A 20 N object displaces 2,000 ml of water. What is the buoyancy? The answer is 19.6 N. Can someone tell me how it got to be this number?

OpenStudy (anonymous):

What are the units of 19.6??

OpenStudy (anonymous):

oh sorry--it's Newtons.

OpenStudy (anonymous):

Fb = rho x g x V where Fb is Force of buoyancy, rho is density of the water, g is acceleration of gravity, and V is volume of the object. rho = 1000 (for density of the water) g = 9.8 (acceleration of gravity) V = 2000 ml = 2 l = 2 dm^3 = 2 x 10 ^-3 m^3 (for volume of the object) Fb = rho x g x V Fb = 1000 x 9.8 x 2 x 10 ^-3 Fb = 9.8 x 2 Fb = 19.6 Newton

OpenStudy (anonymous):

wait...how are you calculating the volume? and where did 1000 for the density come from? :(

OpenStudy (anonymous):

for volume: do you know that 1 L is equal to 1 dm^3??? for density of the water is usually used for the value of 1000, obtained by experimental results, You can get this value from the backyard of a physics textbook or on wikipedia

OpenStudy (anonymous):

sorry if I'm difficult. I'm just having trouble grasping physics. I'm remembering the density of water now but I thought it was 1.0 g...and what is "dm" meant to stand for?

OpenStudy (anonymous):

yes in another unit the density of the water is 1 gram/cm^3.., But if we do the conversion in units of kg/m ^ 3: 1 gram/cm^3 = (1 x10^-3 kg) / (1 x 10^-6 m^3) = 1x 10^3 kg/m^3 = 1000 kg/m^3

OpenStudy (anonymous):

Sorry I forgot to tell you if the unit is for 1000 kg/m ^ 3

OpenStudy (anonymous):

@craftshark98

OpenStudy (anonymous):

thank you...I'm still trying to grasp this

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

sorry i forgot to tell you that "dm" is Decimetre